Circular No. 6071

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

POSSIBLE GRAVITATIONAL MICROLENSING EVENT S. Benetti, L. Pasquini, and R. West, European Southern Observatory

(ESO), report: “Spectra of this object (cf. IAUC 6068, 6069) have now been obtained with telescopes at La Silla during five consecutive nights (Sept. 2–6). No spectral changes are evident in the range 400–700 nm, lending support to the interpretation of this event as being caused by a genuine MACHO (Massive Compact Halo Object). A provisional, quantitative analysis (based on criteria in the range 400–450 nm; West 1972, Bull. Abast. Obs. 43, 109) of a high-S/N 15-min spectrum (resolution 1.7 nm), obtained with the ESO 3.6-m telescope (+ EFOSC2) on Sept. 5.1 UT, yields spectral class K0 IV and M(V) = +3; this indicates that the star is likely to be somewhat reddened and located in the galactic bulge.”

SUPERNOVA 1994X IN ANONYMOUS GALAXY A. V. Filippenko, A. J. Barth, and T. Matheson, University of Califor-

nia at Berkeley; and J. A. Burrous, Lick Observatory, report that preliminary inspection of CCD spectra (range 310–1000 nm) obtained on Sept. 1 UT with the Shane 3-m reflector at Lick suggests that this object is of type Ia, a few months past maximum brightness. The redshift of the parent galaxy is 0.058, based on measurements of the Ca II H+K absorption lines.

PERIODIC COMET MACHHOLZ 2 (1994o) P. Pravec, Ondřejov, reports his discovery of fourth and fifth com-

ponents to this comet on CCD images obtained with the 0.65-m reflector on Sept. 4.1 UT. The fourth component — also independently reported by W. Johnson (Anza, CA), by T. Puckett, J. Armstrong, and M. Marcus (Atlanta, GA), and by M. Jäger (Vienna, Austria) — was then similar in brightness to the third component (IAUC 6070), about a magnitude fainter than the second component (IAUC 6066), sporting a 2′.0 coma with little condensation, and located some 320′′ north-northeast of the primary component. The fifth fragment (also evidently found by Jäger on his films) was nearly 1 mag fainter than the fourth component, consisted of a 1′–2′ diffuse coma and little or no condensation, and was located 307′′ north-northeast of the second component.

1994 September 6 Daniel W. E. Green